### v1

# What You See Is What Matters: A Novel Visual and Physics-Based Metric for Evaluating Video Generation Quality

Abstract
As video generation models advance rapidly, assessing the quality of generated videos has become increasingly critical. Existing metrics, such as Fréchet Video Distance (FVD), Inception Score (IS), and ClipSim, measure quality primarily in latent space rather than from a human visual perspective, often overlooking key aspects like appearance and motion consistency to physical laws. In this paper, we propose a novel metric, VAMP (Visual Appearance and Motion Plausibility), that evaluates both the visual appearance and physical plausibility of generated videos. VAMP is composed of two main components: an appearance score, which assesses color, shape, and texture consistency across frames, and a motion score, which evaluates the realism of object movements. We validate VAMP through two experiments: corrupted video evaluation and generated video evaluation. In the corrupted video evaluation, we introduce various types of corruptions into real videos and measure the correlation between corruption severity and VAMP scores. In the generated video evaluation, we use state-of-the-art models to generate videos from carefully designed prompts and compare VAMP’s performance to human evaluators’ rankings. Our results demonstrate that VAMP effectively captures both visual fidelity and temporal consistency, offering a more comprehensive evaluation of video quality than traditional methods.
11footnotetext: Equal contribution.22footnotetext: Project lead and corresponding author.
1 Introduction
Figure 1: Comparison of Embedding-Based Metrics and Visual Physics-Based Metrics. The figure illustrates the evaluation pipelines for generated video quality. The top section represents embedding-based metrics, which rely on extracting features from generated videos and comparing them in a latent space against reference videos. The bottom section introduces a human-visual-system-inspired (HVS) evaluation framework, which analyzes videos based on visual appearance (color, shape, texture) and motion coherence (realistic movement dynamics).
Recent advancements in video generation, particularly through diffusion models, have dramatically increased the potential for producing highly realistic and dynamic video content [13, 11]. These models enable sophisticated video synthesis, opening up applications in movie making, virtual reality, autonomous systems, and robotics [18, 2, 6, 30, 27, 28]. However, despite these advancements, evaluating the quality of generated videos remains challenging. Existing metrics, such as Fréchet Video Distance (FVD) [24], Inception Score (IS) [21] and ClipSim [19], offer only partial insights into video quality and suffer from significant limitations. For instance, FVD assesses statistical similarity between real and generated videos but lacks interpretability, making its scores difficult to intuitively understand in terms of video quality, while ClipSim captures high-level semantic alignment but can miss finer details, such as texture quality, minor artifacts, or slight color mismatches, which could affect human perception of quality but not be reflected in the score. As video generation models advance in both complexity and realism, these limitations become increasingly problematic. To truly evaluate the quality of generated videos, a comprehensive metric should go beyond latent representations and instead align with standards rooted in human visual perception. Human perception encompasses both visual appearance—covering aspects like color, texture, and shape consistency—and motion coherence, ensuring that movement appears realistic and adheres to natural dynamics. This motivates the development of a new evaluation framework that incorporates both the visual and physical dimensions of video quality, capturing the nuanced standards by which humans assess video realism [31].
Evaluating the quality of generated videos introduces a set of unique and complex challenges, distinct from the assessment of static images. Unlike images, videos require not only high visual fidelity but also temporal consistency, ensuring coherent visual appearance and with and smooth transitions across frames [23]. Capturing this temporal consistency and coherence is essential, as even minor inconsistencies in visual appearance and movement can break the realism of generated videos. Furthermore, generated videos must adhere to the laws of physics, maintaining geometric consistency within scenes and ensuring that objects and movements behave in a physically plausible manner [32]. Beyond physical coherence, semantic consistency is crucial—the content and actions depicted in a video must be logically meaningful and align with real-world expectations [16]. This involves not only assessing visual quality but also ensuring that the video’s narrative or message makes sense. In some cases, particularly when audio is integrated, the evaluation process must also account for multimodal coherence, considering both visual and auditory elements where applicable [34]. These interconnected challenges—spanning visual, physical, and semantic dimensions—highlight the difficulty of developing a truly comprehensive evaluation metric for generated videos.
To address these challenges, we propose a novel metric designed specifically to evaluate generated video quality, with a focus on both visual and motion consistencies. Grounded in a human visual perspective, this metric offers a more intuitive and comprehensive measure of video quality. Our contributions are as follows:
  • We introduce a new video quality evaluation metric that directly targets both visual appearance consistency and motion consistency in generated videos. By moving beyond the limitations of existing metrics like FVD and CLIP score, our approach emphasizes temporal visual appearance and motion realism to provide a more accurate reflection of video quality.
  • We create a corrupted video dataset which includes five types of corruption: brightness, Gaussian noise, impulse noise, black box, and defocus blur. Each type of corruption is applied at five distinct levels of severity, allowing for a granular analysis of how different types and intensities of distortions impact video quality assessment for our and other metrics.
  • We conduct extensive experiments on both real and synthetic video datasets, applying our metric to state-of-the-art video generation models. This analysis provides an in-depth comparison of different metrics to measure the quality of the videos, offering insights into their performance on various types of video data. We show that our proposed metric provides a nuanced and more accurate assessment of generated video quality compared to existing approaches.
By addressing these key dimensions—visual appearance and motion consistencies—our work offers a valuable tool to measure the quality of video generation. We expect that this metric will provide a robust foundation for future advancements of video generation, supporting researchers in the pursuit of more realistic and coherent generated video content [15].
2 Related Work
Image Generation The field of image generation has seen remarkable progress in recent years, primarily driven by advances in deep learning techniques. Generative Adversarial Networks (GANs) [8] have been at the forefront of this revolution, with architectures like StyleGAN [12] and BigGAN [1] pushing the boundaries of image quality and diversity. More recently, diffusion models [10] have emerged as a powerful alternative, demonstrating exceptional image quality and flexibility in tasks such as text-to-image generation [20][4]. These advancements in image generation have laid the groundwork for tackling the more complex task of video generation, providing crucial insights into high-dimensional data synthesis and the importance of perceptual quality in generated content.
Video Generation Building upon the success of image generation models, video generation has become an active area of research, presenting unique challenges due to the temporal dimension. Early approaches extended GANs to the video domain, with models like VGAN [25] and MoCoGAN [23] attempting to capture both motion and content. Recent work has explored the use of 3D convolutions and attention mechanisms, as seen in Video Transformer Network [16], to better model long-range dependencies in video sequences. The advent of large language models has also influenced video generation, with text-to-video models like Make-A-Video [15] demonstrating impressive results in generating videos from textual descriptions. Exploration on using diffusion models led to improvement on text-guided video generation quality. [33]. Despite these advancements, consistent long-term motion and adherence to physical laws remain significant challenges in the field of video generation.
Metrics for Video Generation As video generation models have evolved, so too has the need for robust evaluation metrics. Traditional image quality metrics like Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) have been adapted for video [29], but they often fail to capture the complexities of human perception in video quality assessment. The Fréchet Video Distance (FVD) [24], an extension of the Fréchet Inception Distance (FID) used in image generation, has become a popular metric for evaluating the quality and diversity of generated videos. More recently, perceptual metrics leveraging large pre-trained models, such as CLIP score [19], have been applied to video evaluation, offering a more semantic assessment of generated content. Leveraging multi-modal language model and large-scale human feedback to develop evaler demonstrate another promising direction. [9] However, these metrics still struggle to fully capture aspects such as temporal consistency, physical plausibility, and long-term coherence in generated videos. Our work aims to address these limitations by introducing a novel metric that specifically targets these crucial aspects of video quality.
3 VAMP: Visual Appearance and Motion Plausibility Metric for Video Quality Assessment
To effectively assess the quality of generated videos, it requires metrics that can capture both visual fidelity and motion coherence, aligning with human perception. We introduce VAMP, a novel metric designed to address both visual appearance and motion consistencies of generated videos. VAMP has three main steps: point sampling, mask selection and VAMP score construction.
3.1 Point Sampling and Mask Selection
To properly track and identify the region of interest in the image to evaluate, we use the above mentioned point sampling methods to identify the important regions to track. We then use SAM2 to track the movement of those important regions. We mainly experimented with two selection methods: 1) SAM2-based sampling and 2) SIFT-based Sampling, which have trade-off between quality and efficiency.
One natural way of identifying the regions of interest is utilizing SAM2’s auto segmentation ability where SAM2 identify key features in the initial images. We then use the centroids of the key regions as input points to initiate the tracking. The segmentation masks of the regions of interest are then used for downstream calculations.
An alternate to SAM2 is to use light-weight image features to identify regions of interest. We propose using a combination of SIFT and DBSCAN clustering [5]. DBSCAN groups packed SIFT features together and drop outliers. It allows us to automatically merge detected points together forming variable numbers of regions instead of pre-defined numbers of regions. As this is a CPU-only process, we can alleviate the compute burden on GPU. We show the results of this method in the supplement material.
3.2 VAMP Score
Figure 2: Pipeline for the VAMP Score Calculation. This diagram illustrates the process for calculating the VAMP score. Starting with video input, point sampling and SAM masking are applied to identify and track objects across frames. The pipeline computes the Appearance Score, incorporating metrics for color similarity , shape similarity, and texture similarity. Concurrently, the Motion Score is calculated by evaluating velocity and acceleration consistency, capturing the physical plausibility of object movements. These two scores are combined using weighted factors to compute the final VAMP score.
Our proposed metric for evaluating generated video quality consists of two main components: an appearance score and a motion score. Together, these components offer a comprehensive evaluation of both the visual appearance consistency and the motion consistency of generated video sequences. The appearance score captures visual consistency in terms of color, shape, and texture across frames, while the motion score assesses the smoothness of movements of individual objects across frames.
3.2.1 Appearance Score
Among studies of the human visual system, color, shape, and texture are three fundamental attributes that play a critical role in visual perception [7]. These attributes form the basis for how humans interpret and recognize objects and scenes in their environment. Color provides information about material properties and lighting conditions, shape conveys structural and spatial information, and texture offers fine-grained details about object surface characteristics. Together, they enable humans to perceive, differentiate, and understand objects and surrounding environments with accuracy, making them essential components for evaluating the visual fidelity of generated videos.
Therefore, we design an appearance score to measure the visual consistency of objects across frames. The appearance score consists of three sub-scores that assess color similarity, shape similarity, and texture similarity for each detected objects across all frames of the video. These sub-scores work together to ensure that the objects in frames not only appear realistic individually but also maintain visual coherence when viewed in sequence.
Color Similarity. Color similarity is evaluated by comparing the color distribution between consecutive frames. To quantify this, we employ the Earth Mover’s Distance (EMD), which is also known as Wasserstein distance, to measure the dissimilarity between the color distributions between corresponding objects:
$\text{EMD}(O_{i,t},O_{i,t+1})=\min_{\mathbf{f}}\sum_{j,k}f_{j,k}d_{j,k},$ (1)
where $O_{i,t}$ and $O_{i,t+1}$ are the normalized color histograms of the same corresponding object, $O_{i}$, across frames at time $t$ and $t+1$, $f_{j,k}$ represents the flow between bins $j$ and $k$, and $d_{j,k}$ denotes the ground distance between histogram bins. Lower EMD values indicate greater similarity between the color distributions of objects in consecutive frames. EMD is particularly suited for measuring color similarity because it accounts for the spatial arrangement of color distributions, which ensures that subtle shifts in color distributions can be captured. Since there are three channels of the object image, color similarity score is
$S_{color}(O_{i,t},O_{i,t+1})=1-\frac{1}{3}\sum_{c=\{r,g,b\}}\text{EMD}(O_{i,t,c},O_{i,t+1,c})$ (2)
By using EMD, the color similarity score captures the perceptual alignment of colors, ensuring that generated videos maintain realistic and consistent color fidelity across frames.
Shape Similarity. Shape similarity ensures that objects within consecutive frames maintain their structural integrity and geometry. To evaluate shape similarity, we use the Hausdorff distance, a metric for comparing the geometric consistency of object contours. This measure captures both global and fine-grained shape differences, ensuring that the structural properties of objects remain consistent across frames.
The directed Hausdorff distance between two point sets of corresponding object $O_{i}$ at time $t$, $O_{i,t}$, and object at time $t+1$, $O_{i,t+1}$ is defined as:
$h(O_{i,t},O_{i,t+1})=\max_{o_{i,t}\in O_{i,t}}\min_{o_{i,t+1}\in O_{i,t+1}}\|o_{i,t}-o_{i,t+1}\|,$ (3)
[…]
The final shape similarity score is computed using the symmetric Hausdorff distance, which considers the maximum of the directed distances in both directions:
$S_{\text{Hausdorff}}=\frac{1}{1+\max(h(O_{i,t},O_{i,t+1}),h(O_{i,t+1},O_{i,t}))},$ (4)
and the shape similarity score can be written as:
$S_{shape}(O_{i,t},O_{i,t+1})=\frac{1}{1+S_{\text{Hausdorff}}(O_{i,t},O_{i,t+1})}.$ (5)
By using the Hausdorff distance, the shape similarity score effectively captures the geometric alignment of object contours across frames, ensuring structural consistency throughout the video.
Texture Similarity. Texture similarity assesses the fine-grained surface details of objects across consecutive frames, ensuring that surface patterns remain consistent throughout the video. To evaluate texture similarity, we resize the images of corresponding objects to a common target size and convert them to grayscale for uniform processing. We then compute texture features using the Gray Level Co-occurrence Matrix (GLCM), which captures spatial relationships between pixel intensities, a key indicator of texture.
The GLCM is calculated for multiple orientations (00, $\pi/4$, $\pi/2$, and $43\pi/$ 4) and a distance of 1 pixel, with intensity levels quantized to 256 for consistency. From the GLCM, we extract five statistical properties—contrast, dissimilarity, homogeneity, energy, and correlation—that comprehensively describe texture. These properties are flattened into feature vectors for each object in consecutive frames.
The similarity between the texture features of two frames is then computed using cosine similarity:
$S_{\text{texture}}(O_{i,t},O_{i,t+1})=\frac{\mathbf{f}_{O_{i,t}}\cdot\mathbf{f}_{O_{i,t+1}}}{\|\mathbf{f}_{O_{i,t}}\|\|\mathbf{f}_{O_{i,t+1}}\|},$ (6)
where $\mathbf{f}_{O_{i,t}}$ and $\mathbf{f}_{O_{i,t+1}}$ are the feature vectors derived from the GLCM properties of the corresponding objects at time $t$ and $t+1$. Higher cosine similarity values indicate greater alignment of texture details between frames. This approach ensures that texture consistency can be accurately captured.
Appearance Score. The overall appearance score evaluates the visual fidelity and consistency of objects across frames by integrating three key components: color similarity, shape similarity, and texture similarity. These components collectively ensure that the generated video maintains coherence in color distribution, geometric structure, and fine-grained surface details throughout the sequence.
The final appearance score is computed as the weighted average of the individual similarity scores:
$S_{\text{appearance}}=w_{\text{c}}S_{\text{color}}+w_{\text{s}}S_{\text{shape}}+w_{\text{t}}S_{\text{texture}},$ (7)
where $S_{\text{color}},S_{\text{shape}},\text{and }S_{\text{texture}}$ are the respective similarity scores for color, shape, and texture, and $w_{\text{c}},w_{\text{s}},\text{and }w_{\text{t}}$ are their corresponding weights, normalized such that:
$w_{\text{color}}+w_{\text{shape}}+w_{\text{texture}}=1.$ (8)
These weights can be tuned based on the importance of each component in the specific application. By combining these measures, the overall appearance score provides a comprehensive evaluation of the visual quality of generated videos, aligning with human perception of visual coherence.
3.2.2 Motion Score
The motion score evaluates the physical plausibility of object movements in the generated video. It incorporates several components designed to capture the realism of motion, including velocity consistency and acceleration consistency. These components ensure that movements adhere to natural laws and that the generated video maintains temporal coherence.
Velocity Consistency. Velocity consistency evaluates the stability of object motion by analyzing the trajectories of object centroids across frames. For each object, the velocities are calculated as the differences between the centroids of consecutive frames. The average distance between these centroids is measured using:
$v_{t}=\|\mathbf{c}_{t+1}-\mathbf{c}_{t}\|,$ (9)
where $\mathbf{c}_{t}$ and $\mathbf{c}_{t+1}$ are the centroids of the object in frames $t$ and $t+1$. To quantify the consistency of velocity, we compute the standard deviation of the velocity values across all frames and normalize it by the mean velocity. The velocity consistency score is then defined as:
$S_{\text{vel}}=\exp\left(-\frac{\text{std}(\mathbf{v})}{\text{mean}(\mathbf{v})}\right),$ (10)
Lower velocity variance indicates smoother and more consistent motion.
Acceleration Consistency. Acceleration consistency evaluates the smoothness of changes in velocity across frames. For each object, accelerations are computed as the differences between velocities in consecutive frames:
$a_{t}=v_{t+1}-v_{t},$ (11)
where $v_{t}$ and $v_{t+1}$ represent the velocities between frames $t$ and $t+1$. The variance of accelerations is then calculated, as lower acceleration variance indicates smoother motion transitions. The acceleration consistency score is defined as:
$S_{\text{acc}}=\exp(-\text{Var}(\mathbf{a})),$ (12)
where $\mathbf{a}$ is the vector of accelerations across all frames for a given object.
Motion Score. The final motion score for each object is computed as the weighted combination of the individual components:
$S_{\text{motion}}=w_{\text{vel}}S_{\text{vel}}+w_{\text{acc}}S_{\text{acc}},$ (13)
where $w_{\text{vel}},w_{\text{acc}},\text{and }w_{\text{depth}}$ are the weights for velocity, acceleration, and depth consistency, respectively, satisfying:
$w_{\text{vel}}+w_{\text{acc}}=1.$ (14)
By integrating these components, the motion score provides a comprehensive assessment of the physical plausibility and temporal coherence of object movements in generated videos.
Figure 3: Evaluation of video quality using the VAMP metric across different levels of corruptions and generative models. This figure shows the impact of various corruption types on video quality as assessed by the VAMP metric. (Left) Examples of corrupted videos showing the effect of Brightness, Gaussian Noise, Impulse Noise, Defocus Blur, and Black Shapes at different severity levels. (Right) Generated video outputs from state-of-the-art video generation models (T2VZ, Pika, MS, VC2) based on the given prompt.
3.2.3 VAMP Score
By combining the visual appearance and motion scores, our proposed metric provides a comprehensive evaluation of video generation quality, ensuring that both visual fidelity and physical plausibility are considered. This holistic approach allows for more accurate and intuitive assessments of generated video content.
The VAMP score of a video is computed as a weighted sum of the appearance score ($S_{\text{appearance}}$) and the motion score ($S_{\text{motion}}$) across all the frames and objects:
$S_{\text{VAMP}}=\frac{1}{mn}\sum_{t=0}^{n}\sum_{i=0}^{m}\alpha S_{\text{appearance},i,(t,t+1)}+\beta S_{\text{motion},i,(t,t+1)}$ (15)
where $m$ is the total number of objects, $n$ is total number of frames, $\alpha$ and $\beta$ are the weights assigned to the appearance and motion scores, respectively, and satisfy the normalization condition:
$\alpha+\beta=1.$ (16)
These weights allow the metric to be adjusted based on the specific requirements of the application. For instance, applications where visual fidelity is more critical may assign a higher value to $\alpha$, while scenarios requiring precise motion realism may prioritize $\beta$.
Interpretability. A higher $S_{\text{VAMP}}$ score reflects better alignment with human perception of video quality, encompassing both visual appearance and object movement aspects. This makes $S_{\text{VAMP}}$ an intuitive and practical metric for evaluating the quality of generated video content across diverse domains. By combining these complementary dimensions, the VAMP score bridges the gap between traditional evaluation metrics and the nuanced expectations of human observers, offering a robust foundation for benchmarking video generation models.
4 Experiments
Method Brightness Gaussian Noise
0 1 2 3 4 5 0 1 2 3 4 5
ClipSim ↑ 0.969 0.968 0.967 0.966 0.965 0.965 0.969 0.948 0.944 0.945 0.945 0.944
IS ↑ 1.470 1.494 1.506 1.523 1.533 1.537 1.470 1.555 1.628 1.721 1.829 1.816
FVD (×$\times 10^{3}$) ↓ 0 0.239 0.657 1.121 1.721 2.429 0 1.759 2.974 4.295 5.395 6.314
VAMP-A ↑ 0.726 0.729 0.713 0.727 0.721 0.712 0.726 0.694 0.661 0.663 0.562 0.541
VAMP-M ↑ 0.645 0.645 0.632 0.624 0.613 0.607 0.645 0.604 0.565 0.519 0.439 0.388
VAMP ↑ 0.685 0.670 0.656 0.655 0.646 0.638 0.685 0.631 0.594 0.562 0.476 0.434
Impulse Noise Defocus Blur Black Shapes
0 1 2 3 4 5 0 1 2 3 4 5 0 1 2 3 4 5
0.969 0.948 0.946 0.946 0.943 0.943 0.969 0.970 0.970 0.972 0.973 0.975 0.969 0.962 0.955 0.951 0.948 0.944
1.470 1.679 1.728 1.775 1.876 1.846 1.470 1.418 1.421 1.419 1.395 1.367 1.470 1.496 1.509 1.539 1.561 1.596
0 3.515 4.350 4.885 5.789 6.431 0 1.244 1.949 3.355 4.575 5.290 0 2.362 1.722 1.578 1.866 2.230
0.726 0.726 0.711 0.699 0.660 0.582 0.726 0.721 0.698 0.580 0.372 0.220 0.726 0.683 0.681 0.659 0.647 0.642
0.667 0.621 0.594 0.566 0.501 0.404 0.667 0.640 0.605 0.477 0.281 0.154 0.667 0.636 0.573 0.533 0.523 0.519
0.685 0.653 0.629 0.606 0.549 0.458 0.685 0.664 0.633 0.508 0.308 0.173 0.685 0.650 0.606 0.571 0.560 0.556
Table 1: Performance of Metrics Across Corruption Types and Levels. This table shows the evaluation of video quality metrics under different corruption types (Brightness, Gaussian Noise, Impulse Noise, Defocus Blur, and Black Shapes) applied at levels 1–5. Metrics include ClipSim, IS, FVD (×$\times 10^{3}$), VAMP-A (appearance-only), VAMP-M (motion-only), and VAMP.
We evaluate the robustness and effectiveness of our proposed metric in two scenarios: corrupted video evaluation and generated video evaluation. To benchmark its performance, we compare it against established metrics such as Fréchet Video Distance (FVD), Inception Score (IS), and ClipSim. The following experiments demonstrate how our metric captures both visual fidelity and temporal consistency in videos and outperforms baseline methods.
4.1 Corrupted Video Evaluation
In this experiment, we aim to evaluate the sensitivity of our proposed metric by injecting various types of noise and corruption into real videos.
Dataset. Given the wide range of possible video scenes and objects, it is impractical to evaluate all types of video content. In this study, we focus on videos involving humans with actions, which are among the most common types of objects generated by video generation models. We use the UCF101 dataset [22], which contains a diverse collection of human actions across 101 categories within different scenes, making it ideal for evaluating videos involving human appearance and motion.
Corruption Types. In this experiments, we apply various types of noise to videos from each dataset, including brightness, gaussian noise,impulse noise, defocus blur, and black shapes. Our goal is to measure how well the metric captures the impact of these corruptions across different objects and scene types. The correlation between the noise severity and the metric score will reveal the robustness and effectiveness of the metric in detecting quality degradation in various real-world video contexts. We apply a variety of common image corruption techniques to real videos, simulating different types of noise and distortions. Below is a brief description of each type of corruption used in the experiments :
  • Brightness: Modifies the overall brightness of the video frames, simulating changes in lighting conditions.
  • Gaussian Noise: Adds Gaussian noise to the video frames, simulating random variations in brightness or color. This type of noise is often used to simulate sensor noise.
  • Impulse Noise (Salt-and-Pepper Noise): Adds random white and black pixels to the video, simulating the effect of faulty pixel readings in a sensor or transmission errors.
  • Defocus Blur: Simulates the effect of the video being out of focus, causing a uniform blur across the frames.
  • Black Shapes: Adds random black boxes to the video frames, simulating occlusions, artifacts, or sensor malfunctions.
Experiment Setup. In this experiment, we applied noise to a set of real videos at varying levels (1-5) to simulate different degrees of corruption. Each type of corruption was systematically injected into the original videos, altering the visual quality to different extents. After introducing these distortions, we used our proposed metric to evaluate the resulting videos, generating both an appearance score and a motion score for each, which are used to compute VAMP score. Then, we compare the score calculated from our metric with the baseline methods to show the effectiveness on reflecting the correlations between corruption levels and scores.
4.2 Generated Video Evaluation
In addition to evaluating our metric on corrupted real videos, we further assess its effectiveness by applying it to videos generated by current state-of-the-art video generation models.
Experiment Setup. For this evaluation, we sampled 3,000 prompts from the VidProM dataset, which includes a diverse range of video prompts and scenarios. Four video generation models were used to produce these videos: Pika [17], VideoCraft2 [3] , Text2Video-Zero [14] , and ModelScope [26] . The sampled prompts were used by the video generation models to generate corresponding videos. With the generated videos, we evaluate them using our proposed metric, calculating both the appearance score and motion score for each video. The evaluation provides insights into how well the generated videos maintain visual fidelity and motion coherence across different models and scenarios.
5 Results and Analysis
Figure 4: Normalized changes in metric values across corruption types and levels relative to the original video. Each heatmap represents the impact of a specific corruption type on four evaluation metrics (ClipSim, IS, FVD, and VAMP). Blue indicates a positive correlation, while red indicates a negative correlation.
Results. The evaluation results on corrupted videos are summarized in Table 1, which presents the performance of video quality metrics under different corruption types (Brightness, Gaussian Noise, Impulse Noise, Defocus Blur, and Black Shapes) applied at varying levels (1–5). The table includes scores for ClipSim, IS, FVD, VAMP-A (appearance-only), VAMP-M (motion-only), and VAMP. Table 2 reports the scores of generative models (Text2Video-Zero, ModelScope, VideoCrafter2, and Pika) across appearance metrics (Color, Shape, Texture), motion score, and the overall VAMP score. Figure 4 illustrates the normalized changes in metric values for each corruption type and level relative to the original video, showing the effect of corruption on metrics such as ClipSim, IS, FVD, and VAMP. Based on these results, we highlight some key findings in the following paragraphs regarding the sensitivity of the metrics to different corruption types and the generative models’ performance across various metrics.
VAMP captures both visual appearance and motion consistencies of videos. The results presented in Table 1 and Table 2 demonstrate that the VAMP metric effectively evaluates both appearance and motion aspects of video quality. For appearance, VAMP-A specifically quantifies attributes such as color, shape, and texture, as evidenced by its sensitivity to degradation levels in Table 1. For motion consistency, VAMP-M highlights temporal coherence across frames, maintaining robust scores despite challenging corruptions like Impulse Noise and Defocus Blur. Furthermore, Figure 4 illustrates that VAMP balances responses to visual and motion degradations, showing incremental changes across corruption levels that align with human perceptual degradation. These results support VAMP’s capability to holistically assess video quality by incorporating both spatial and temporal dimensions.
VAMP can be a good reference-free alternative to FVD. The analysis of Table 1 and Figure 4 demonstrates that the VAMP score performs comparably to FVD across diverse corruption types and levels. Both metrics reliably capture increasing degradation in video quality as corruption levels intensify, with FVD generally exhibiting sharper increases, particularly under Gaussian Noise and Impulse Noise. However, a key advantage of VAMP is its independence from reference videos, unlike FVD, which requires access to ground-truth data for effective evaluation. This feature makes VAMP more versatile and applicable in real-world scenarios where reference videos may not always be available or practical to use. Figure 4 further highlights importance of VAMP’s evaluation on visual appearance not in latent space. This is particularly evident in corruption types such as Black Shapes, where VAMP exhibits greater stability and consistency in its response compared to FVD, which shows negative correlations to corruption levels. Overall, the VAMP metric emerges as a robust and interpretable alternative to FVD, particularly suited for use cases where reference videos are not accessible.
ClipSim and IS may not be reliable metrics for evaluating generated video quality in certain cases. As shown in Table 1, IS exhibits only minor changes across different corruption levels for Brightness and Gaussian Noise, suggesting it is less sensitive to perceptual degradations. Similarly, ClipSim remains largely unchanged across all tested corruption types, as seen in Figure 4, with negligible variations even for severe distortions like Impulse Noise and Defocus Blur. This insensitivity implies that neither IS nor ClipSim adequately reflects the impact of degradations on visual quality or motion consistency. For example, while Black Shapes significantly affect the motion and appearance of videos, ClipSim fails to capture these changes effectively, maintaining nearly constant values. These limitations highlight that IS and ClipSim might not fully represent the quality of generated videos, particularly in scenarios where both visual appearance and temporal consistency are critical for evaluation.
Model Color Shape Texture Motion VAMP
Text2Video-Zero - - - 0.448 0.314
ModelScope 0.909 0.414 0.956 0.631 0.650
VideoCrafter2 0.957 0.448 0.975 0.683 0.696
Pika 0.993 0.477 0.997 0.788 0.778
Table 2: Quantitative evaluation of generative video models across appearance and motion metrics. The table presents the scores for four video generation models on appearance metrics (Color, Shape, and Texture), Motion Score, and the overall VAMP Score.
6 Conclusion
In this paper, we introduced VAMP, a novel metric for evaluating both visual fidelity and motion coherence in generated videos. By assessing visual appearance consistency and the motion consistency, VAMP provides a more comprehensive evaluation than traditional metrics like IS, and ClipSim, and achieved compatible results as FVD without the need of reference videos. Our experiments show that VAMP effectively captures video quality degradation and aligns well with human evaluations on visual appearance and motion consistency.
Limitations and Future Work. VAMP involves extensive parameter tuning which can be time-consuming and may not generalize well. Its computational complexity presents challenges due to resource-intensive object segmentation and tracking from SAM. Future work can explore adaptive tuning mechanisms and optimize the computational pipeline to enhance efficiency and scalability.